OpenStudy (bananas1234):

Use the quadratic formula to solve x^2 + 9x + 10 = 0. Estimate irrational solutions to the nearest tenth. Options. {–3.6, 16.4} {–7.7, –1.3} {–15.4, –2.6} {–25, 16}

OpenStudy (anonymous):

try pluging in the points

OpenStudy (anonymous):

you should get the awnser

OpenStudy (bananas1234):

@rational

OpenStudy (bananas1234):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

\[\Large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] \[\Large x=\frac{-9 \pm \sqrt{9^2-4*1*10}}{2*1}\] \[\Large x=\frac{-9 \pm \sqrt{41}}{2}\] I'll let you finish up

OpenStudy (bananas1234):

option c?

OpenStudy (whpalmer4):

Actually, this equation is easy enough to solve in your head to find the answer choice! You can also check your answer (it is incorrect) by plugging into the formula. if \(x=-2.6\) as your answer choice suggests, then \[x^2+9x+10=0\]or\[(-2.6)(-2.6)+9(-2.6)+10=0\]\[6.76-23.4+10=0\]Uh, no, that's not true, so that is not a correct answer.

OpenStudy (whpalmer4):

Jim worked the entire problem for you, except evaluating it to a decimal. Can you tell me how you came up with option c?

OpenStudy (bananas1234):

becuase 

OpenStudy (whpalmer4):

While that is true, that's not the entire formula he gave you — you have to divide that by 2.

OpenStudy (whpalmer4):

Option C is what I like to call a sucker answer — looks like a value you'll get at an intermediate point in the computation, or a value you'll get if you make a frequent mistake. You see something on your paper or calculator that matches one of the answers and say "that must be it!" without bothering to check if it actually fits the problem.

OpenStudy (bananas1234):

oh, i see so it must be option B.

OpenStudy (bananas1234):

Thank you for your help

OpenStudy (bananas1234):

:)

OpenStudy (whpalmer4):

By the way, you can work this one out pretty easily in your head. \[x^2+9x+10=0\]For the moment, let's pretend that "\(+10\)" isn't there. \[x^2+9x\cancel{+10}=0\]\[x^2+9x=0\]\[x^2=-9x\]\[\frac{x^2}{x} = \frac{-9x}{x}\]\[x=-9\]So we can guess that a value of \(x\) that satisfies the equation is somewhere in the neighborhood of -9. We can proceed in a number of ways from here to tighten up our estimate. First is to plug that value in for \(x\) and see how close it is: \[(-9)^2 + 9(-9) + 10 = 0\]\[81-81+10=0\]Uh, well, not exactly :-) The \(x^2\) term is positive for all values of \(x\), so we need to make the first term + 10 equal the second term (except the second term will be negative). That means we need the value of \(x^2\) to be about \(10\) less than it is now (81). \(81-10 = 71\) What is a number which multiplied by itself gives about \(71\)? \(9*9 = 81\), how about \(8*8 = 64\)? Let's try \(-8\) in our equation: \[(-8)*(-8) + 9(-8) + 10 = 0\] \[64-72+10 = 0\]\[-8+10=0\]We are getting closer! How about trying \(-7\): \[(-7)*(-7)+9(-7) + 10 = 0\]\[49-63+10 = 0\]\[-14+10=0\]This is not as close as the previous try, so we must have gone too far. But, we now know it is somewhere between \(-7\) and \(-8\), and if we look at our answer choices, there's only one that fits. We plug both of those answer choices into the formula to make sure that they do work and are not a trap.